What is the missing length of a triangle given that one leg is 5 cm long and the other leg is 12 cm long?
Answer by stanbon(75887)

What is the missing length of a triangle given that one leg is 5 cm long and the other leg is 12 cm long?
------
Assuming you have a right triangle,
h^2 = 5^2 + 12^2
h^2 = 25 + 144
h^2 = 169
h = 13 cm
-------------
Cheers,
Stan H.
------------